OpenStudy (daenio):

A single die is rolled. What is the expected die face value, when rolling a single die?

OpenStudy (daenio):

@blockcolder @satellite73

OpenStudy (anonymous):

take the average, i.e. add up the numbers and divide by 6

OpenStudy (daenio):

Is there a specific formula I could use?

OpenStudy (anonymous):

2012Sorching summer comes, in order to thank everyone, characteristic, novel style, varieties, low price and good quality, and the low sale price. Thank everyone ==== ( http://www.fullmalls.com ) ===== ==== ( http://www.fullmalls.com ) ===== $33 True Religion jeans, Ed Hardy jeans,LV,Coogi jeans,Affliction jeans $30 Air Jordan shoes,Shox shoes,Gucci,LV shoes 50%Discount winter fashion :Sandle,t-shirt,caps,jerseys,handbag and brand watches!!! $15 Ed Hardy ,LV ,Gucci Bikini $15 Polo, Ed Hardy, Gucci, LV, Lacoste T-shirts $25 Coach,Gucci,LV,Prada,Juicy,Chanel handbag, $10 Gucci,Ed Hardy sunglasses $9 New Era caps. give you the unexpected harvest ==== ( http://www.fullmalls.com ) ===== ==== ( http://www.fullmalls.com ) ===== ==== ( http://www.fullmalls.com ) ===== ==== ( http://www.fullmalls.com ) ===== ==== ( http://www.scnshop.com ) =====

OpenStudy (daenio):

Is that a site for AUTHENTIC products?

OpenStudy (anonymous):

jesus here we go again with the ads

OpenStudy (daenio):

Alright, thanks @satellite73!

OpenStudy (unklerhaukus):

\[\text{if } x =\text{ the value of the roll,}\]\[\text{ }\langle x\rangle=\text {the expectation value of x,}\]\[N=\text{ is the total number of possibilities} \]\[\text{and the summation is taken over all posible x } \] \[\langle x\rangle=\sum \frac xN\] \[\langle x\rangle=\sum \frac x6=\frac {1+2+3+4+5+6}6=\frac{21}6=3.5\]

OpenStudy (daenio):

I figured it out, but thanks! :D